# On Task Performance and Model Calibration with Supervised and Self-Ensembled In-Context Learning

### v2

@@ 6.1 Comparison between Learning Methods
We first present a comprehensive analysis of various learning methods in low-resource settings, detailing model performance and calibration errors in Table 1. For full experimental results, we refer the readers to Appendix B.
@@
Performance and calibration of learning methods are task-dependent. We find that learning methods perform differently depending on the datasets and we divide the tasks into different families depending on their observed behavior. ICL demonstrates comparable performance to SFT/SICL on SST-2 and RTE. However, tuning on these datasets with SFT/SICL yields increased ECE along with higher NLL and lower IE scores as shown in Table 2, but no substantial in-task performance improvement. This indicates that the model does not recover the ground truth distribution in the test set while becoming more confident and certain about its predictions, which serves as a sign of miscalibration. Conversely, tasks such as intent detection (NLU++), Manifestos, and Hate speech, show noticeable performance enhancement and better calibration with lower ECE by using SFT/SICL. Nevertheless, despite these task-dependent variations, ECE remains relatively high across all methods except for intent detection, indicating the problem of miscalibration across all learning methods.
@@
ICL can achieve comparable performance with SFT on ‘seen’ data. We suspect the divergent behaviors are possibly due to data contamination of FLAN training corpus Longpre et al. (2023) wherein ZSL and ICL exhibit similar performances with SFT and SICL on training datasets labeled as seen (e.g., SST-2, RTE) [^2] We corroborate findings from other concurrent work on data contamination Zhu et al. (2023); Deng et al. (2023) that also reports the unfair practice of evaluations on seen datasets. . To further investigate the performance on seen datasets, we apply the Batch Calibration method Zhou et al. (2023a), as shown in Table 3. Surprisingly, we find that ICL performs on par or even better than SICL with calibration across all the (possibly) seen data, which reveals the ability of LLMs that have been recovered by calibration techniques on these seen tasks. However, for unseen datasets (NLU++, Manifestos, etc.), the performance of ICL is not comparable to either SICL or SFT even with calibration.
  Evaluation Metrics SST-2 SST5
ICL SICL ICL SICL
Performance acc 95.22 95.63 50.48 54.25
macro f1 - - 37.59 47.12
acc 95.95 95.63 50.89 49.91
+ calibrated macro f1 - - 49.80 49.89
Evaluation Metrics Manifestos Hate speech
ICL SICL ICL SICL
Performance micro f1 19.29 38.12 40.18 61.33
macro f1 13.01 37.55 40.09 59.48
micro f1 31.00 38.33 45.11 59.58
+ calibrated macro f1 29.15 37.83 42.36 57.81
Table 3: A selection of results after applying Batch Calibration. We refer the readers to Appendix B.1 for the full set of results with the calibration method.
@@
Different learning methods excel with different task families. Given the comparison of the performances and calibration on different datasets, we suggest that the choice of learning methods should be task-dependent. The experiments and analysis indicate that unseen datasets obtain better performance and more trustworthy results with supervised tuning methods. For the seen datasets, ICL combined with other ‘tweaks’ such as model calibration can be a better choice, since the supervised tuning methods are more likely to make the model over-confident and less trustworthy.
@@
6.2 Self-Ensembling Results
@@
So far, we have observed the common miscalibration issues for all learning methods. We then investigate the feasibility of self-ensembling to improve calibration.
Self-ensembling works across learning methods and enhances calibration performance. In Table 4, with different learning methods combined with self-ensembling variations in designs, we find that by changing the in-context example combinations or prompting templates, the best performance of self-ensembling outperforms the baseline without any ensembling by 0.79. Even though the performance gains seem marginal, self-ensembling substantially enhances the calibration performance, reducing the mean ECE value by 43%. The empirical results also show that in addition to ICL, SFT, and SICL also benefit from self-ensembling in both performance and calibration scores. SFT and SICL exhibit a larger drop in ECE after self-ensembling than ICL. We also notice that when self-ensembling over SFT and SICL, the model has lower ECE scores than ICL, but with much better performances. This indicates the efficiency of self-ensembling in making the predictions more trustworthy while maintaining or even improving the performances. It also indicates that self-ensembling has the potential to mitigate the prominent problem of overconfidence in supervised tuning methods, as shown by Figure 2.
((a)) SICL
((b)) E-SICL(Var-Both)
Figure 2: The confidence histograms and reliability diagrams of SICL and self-ensembling with max probability on SST5. E- means the self-ensembled systems.
Different Variations and Ensembling Strategies. Our results suggest that with ICL, Var-IC yields more improvements than Var-Prompt, while the latter shows its efficiency with SFT and SICL. This difference stems from inappropriate prompting templates, requiring model tuning to adhere to the prompt structure. We also find that combining both variations may not necessarily improve the performance but is helpful in enhancing the trustworthiness empirically. Regarding ensemble strategies, we notice that the majority vote improves the performances but struggles to reduce the calibration error. Ensembling with max probability consistently produces the most faithful predictions with promising performances. This can be explained by the idea that normalizing over the max probabilities in a way smooths the probability distribution, making the max probabilities less extreme and mitigating over-confidence issues.
6.3 Ablation Studies
@@
We explore how varying the number of ensembling components with different in-context examples affects performance and calibration. Figure 3 shows the performance and ECE scores with different components. We observe that although the performances remain comparable, the calibration is improved with more components in both ICL and SICL. This highlights the effectiveness of the self-ensembling method in making the predictions more trustworthy through increased input variations.
Diversity in Var-Prompt influences self-ensembling. In Table 5, we show that when tuning with more templates, SFT has lower calibration errors whereas those of SICL increase. Regarding self-ensembling results, we empirically find that by introducing more prompting templates, self-ensembling yields lower calibration error with SFT, but shows worse calibration with SICL, meanwhile yielding similar performances. Nonetheless, self-ensembling consistently improves the performance and calibration within each setting.
@@
The key findings are robust across different training data sizes. We experiment with larger training data and report the results of SFT and SICL on SST-5 in Figure 4. We observe that both SFT and SICL yield better performance and lower ECE with more training data, whereas ICL maintains similar performance. This indicates that supervised methods can better calibrate the model, and produce more calibrated predictions if provided with sufficient data. We also find that self-ensembling remains effective in improving the performance and mitigating the calibration error of the model on average.
The findings hold across different model sizes. In order to assess the impact of model sizes, we further conduct experiments with Flan-T5xl. For full detailed experiment results, we refer the readers to Table 12 and 13 in the Appendix. We find that a larger model achieves a better calibration score with SFT and SICL than ZSL and ICL. We also witness similar behavior of ICL, surpassing SFT and SICL with or without Batch Calibration on seen data, which aligns with previous findings. The results on SST-5 and Hate Speech show that by applying the self-ensembling method, Flan-T5xl achieves better performance and lower calibration scores, indicating that the model becomes more ‘task-specialized’. It is also worth noticing that our method is able to improve the performance and decrease the calibration scores on some tasks (Hate Speech) where traditional calibration currently fails. In general, this indicates the potential of self-ensembling in improving both task performance and calibration.
@@ 7 Conclusion
We have provided a comprehensive analysis of the intricate relationship between in-task performance and calibration across various learning methods in low-resource scenarios. Our findings illuminate the nuances of in-task performance and calibration across different task families, meanwhile addressing the inherent miscalibration over all learning methods. We have also investigated effective strategies to enhance both aspects simultaneously, offering a viable solution through self-ensembling: it results in more calibrated predictions with comparable or superior task performance. We hope that this study will contribute valuable insights into the dynamic landscape of LLMs. These discoveries also offer practical guidance to practitioners, aiding them in choosing suitable learning paradigms and paving the way for the development of more reliable and high-performing LLMs across diverse applications.
@@
Our experimental results are conducted with the Flan-T5 model family, which is an encoder-decoder architecture, where we have not investigated how the behaviours of other popular choices of decoder-only models, such as Llama Touvron et al. (2023), would behave in low-resource scenarios with different learning methods and self-ensembling strategies. Secondly, limited by the training resources, our experiments only consider LLMs within 3B parameters. We will endeavor to scale our experiments to cover larger language models as part of future work. Moreover, there is a variety of sophisticated ensembling methods Mohammed and Kora (2023), where we have only studied the max, mean, and majority vote variants for self-ensembling. In future work, we aim to extend the analysis and the self-ensembling methods to more families of tasks, diverse types of models, and other ensembling techniques.
@@ B.2 Results of self-ensembling
Appendix C Supplementary Results for Ablations
C.1 How about larger models?

### v3

@@ 5 RQ1: How does ICL Impact Model Calibration?
5.1 Results and Discussions
The language model with ICL is not necessarily calibrated despite of the high in-task performance. We observe that when testing the model performance with the least inductive bias using ZSL, the model shows relatively high ECEs on SST-2 and RTE despite that it achieves more than 80% accuracy, and is only calibrated on SST-5 (ECE <0.2). When we use ICL without tuning the model, we only witness a significant drop of ECE in 2 of 7 datasets. By tuning the model with SFT or SICL, the model becomes even less calibrated on SST-2, RTE, SST-5, although it achieves better in-task performance. Based on this finding, we further look into the interplay between the in-task performance and model calibration.
@@ 5.1 Results and Discussions
Performance and calibration of learning methods are task-dependent, depending on whether the data has been seen by the model. We find that learning methods perform differently depending on the datasets and we divide the tasks into different families depending on their observed behavior. ICL demonstrates comparable performance to SFT/SICL on SST-2 and RTE. However, tuning on these datasets with SFT/SICL yields increased ECE along with higher NLL and lower IE scores as shown in Table 2, but no substantial in-task performance improvement. This indicates that the model does not recover the ground truth distribution in the test set while becoming more confident and certain about its predictions, which serves as a sign of miscalibration. Conversely, tasks such as intent detection (NLU++), Manifestos, and Hate speech, show noticeable performance enhancement and better calibration with lower ECE by using SFT/SICL. Nevertheless, despite these task-dependent variations, ECE remains relatively high across all methods except for intent detection, indicating the problem of miscalibration across all learning methods.
@@
ICL can achieve comparable performance with SFT on ‘seen’ data. We suspect the divergent behaviors are possibly due to data contamination of FLAN training corpus (Longpre et al., 2023) wherein ZSL and ICL exhibit similar performances with SFT and SICL on training datasets labeled as seen (e.g., SST-2, RTE). [^2] We corroborate findings from other concurrent work on data contamination (Zhu et al., 2023; Deng et al., 2023) that also reports the unfair practice of evaluations on seen datasets. To further investigate the performance on seen datasets, we apply the Batch Calibration method (Zhou et al., 2024a), as shown in Table 9 in the Appendix. Surprisingly, we find that ICL performs on par or even better than SICL with calibration across all the (possibly) seen data, which reveals the ability of LMs that have been recovered by calibration techniques on these seen tasks. However, for unseen datasets (NLU++, Manifestos, etc.), the performance of ICL, even with the calibration method applied, is not comparable to those of either SICL or SFT.
5.2 Practical Guideline on Choosing Learning Methods
@@
Given the comparison of the performances and calibration on different datasets, we suggest that the choice of learning methods should be task-dependent. The experiments and analysis indicate that unseen datasets obtain better performance and more trustworthy results with supervised tuning methods. For the seen datasets, ICL combined with other ‘tweaks’ such as model calibration can be a better choice, since the supervised tuning methods are more likely to make the model over-confident and less trustworthy.
@@
6 RQ2: How to Ensure Performance and Calibration?
6.1 Self-Ensembling
So far, we have observed the common miscalibration issues for all learning methods. We then investigate the feasibility of self-ensembling to improve calibration. There are two points that create possible variations that can be used for ensembling: 1) variation in the selection of in-context examples (for ICL and SICL), and 2) variation in the chosen prompt (for all the paradigms). Previous work focuses on 1) selecting a better combination of in-context examples (Su et al., 2023) for the model or 2) generating an optimal prompting template (Zhou et al., 2023). On the other side, how the variation of multiple demonstration combinations and prompting templates influences the model behavior is still unexplored. Furthermore, we can 3) ‘self-ensemble’ the model with different ensembling strategies. We now introduce these variants.
6.1.1 Variation of Ensembling Components
Variation of Prompting Templates (Var-Prompt). Different prompting templates have shown high variance in task performance (Mishra et al., 2022). By changing the wording in the templates $f_{p}$, we can also create variations even with the same input to the model. For each input $x$, we randomly select a prompting template $f_{p}^{\prime}$ from a set of available prompting template candidates. In ICL and SICL, the same template is also applied to the in-context examples, formatting the final input as $[f_{p}^{\prime}(x_{1}),y_{1};...;f_{p}^{\prime}(x_{M}),y_{M},f_{p}^{\prime}(x)]$. This makes it applicable not only to ICL and SICL, but also to ZSL and SFT as well.
6.1.2 Self-Ensembling Strategy
Majority Vote. We select the predicted results that have the highest accumulated probability across $K$ variants as the ensembling predictions. The accumulated probability $\mathcal{P}_{acc}$ for the predicted label $l_{i}$ is defined as $\mathcal{P}_{acc}(\hat{y}=l_{i})=\sum_{k=1}^{K}\mathcal{P}_{k}(\hat{y_{k}}=l_{i})\mathbb{I}(\hat{y_{k}}=l_{i})$. We pick the variants that have the same prediction as the ensembling prediction and average the probability distribution of the selected components $\mathcal{P}_{ens}(y|x)=\frac{1}{K^{\prime}}\sum_{k=1}^{K^{\prime}}\mathcal{P}_{k}(y|x)$ where $K^{\prime}$ is the number of selected variants.
Mean Probability. We average the predicted probability distribution of $K$ variants and use the prediction that has the largest probability in the averaged distribution as the ensemble result $\hat{y_{ens}}={\arg\max}_{j}\ \mathcal{P}_{ens}(y_{j}|x)$, where $\mathcal{P}_{ens}(y|x)$ is described as $\mathcal{P}_{ens}(y|x)=\frac{1}{K}\sum_{k=1}^{K}\mathcal{P}_{k}(y|x)$.
Max Probability. For each possible value in the output space, we find the maximum probability of the predicted values across $K$ variants and use this as the prediction’s probability $\mathcal{P}^{\prime}(\hat{y}=l_{i}|x)=\max(\mathcal{P}_{j}(\hat{y}=l_{i}|x),j\in[1,K])$. Because the probability is obtained from different components, the summation of these probabilities is not guaranteed to be 1. Therefore, we apply the normalization on the new probability distribution: $\mathcal{P}_{ens}(y|x)={\rm Norm}(\mathcal{P}^{\prime}(y|x))$. The ensemble prediction is determined as the $\hat{y}$ that has the highest probability after the ensembling step.
6.2 Results and Discussions
@@
Self-ensembling works across learning methods and enhances calibration performance. In Table 3, with different learning methods combined with self-ensembling variations in the design, we find that by changing the in-context example combinations or prompting templates, the best performance of self-ensembling outperforms the baseline without any ensembling by 0.79. Even though the performance gains seem marginal, self-ensembling substantially enhances the calibration performance, reducing the mean ECE value by 43%. The empirical results also show that in addition to ICL, SFT, and SICL also benefit from self-ensembling in both performance and calibration scores. SFT and SICL exhibit a larger drop in ECE after self-ensembling than ICL. We also notice that when self-ensembling over SFT and SICL, the model has lower ECE scores than ICL, but with much better task performance. This indicates the efficiency of self-ensembling in making the predictions more trustworthy while maintaining or even improving the task performance. It also suggests that self-ensembling has the potential to mitigate the prominent problem of overconfidence in supervised tuning methods, as shown by Figure 1.
Different Variations and Ensembling Strategies. Our results suggest that with ICL, Var-IC yields more improvements than Var-Prompt, while the latter shows its efficacy with SFT and SICL. This difference stems from inappropriate prompting templates, requiring model tuning to adhere to the prompt structure. We also find that combining both variations may not necessarily improve the performance but is helpful in enhancing the trustworthiness empirically. Regarding ensemble strategies, we notice that the majority vote improves the performance in general, but struggles to reduce the calibration error. Ensembling with max probability consistently produces the most faithful predictions with promising performances. This can be explained by the idea that normalizing over the max probabilities in a way smooths the probability distribution, making the max probabilities less extreme and mitigating over-confidence issues.
6.3 Discussions and Ablation Studies
Comparison with Classical Calibration Methods. Compared to Platt and Temperature Scaling Guo et al. (2017), self-ensembling doesn’t require any supervised labelled data to tune the parameters for different instruction templates, ICL settings and models. In addition, classical calibration methods merely adjust confidence levels while preserving the ranking among choices in classification tasks, making it ineffective in enhancing performance (see Table 11 in the Appendix). As a different perspective, self-ensembling is orthogonal to other calibration methods, allowing other calibration methods to be applied on top of the ensembled results.
@@ 6.3 Discussions and Ablation Studies
We explore how varying the number of ensembling components with different in-context examples affects performance and calibration. Figure 2 shows the performance and ECE scores with different components. We observe that although the performances remain comparable, calibration is improved with more components in both ICL and SICL. This highlights the effectiveness of the self-ensembling in making the predictions more trustworthy through increased input variations.
@@
Diversity in Var-Prompt influences self-ensembling. Table 4 show that when tuning with more templates, SFT has lower calibration errors whereas those of SICL increase. Regarding self-ensembling results, we find that by introducing more prompting templates, self-ensembling yields lower calibration error with SFT, but shows worse calibration with SICL, meanwhile yielding similar performances. Nonetheless, self-ensembling consistently improves the performance and calibration in each setting.
The key findings are robust across different training data sizes. We experiment with larger training data and report the results of SFT and SICL on SST-5 in Figure 3. We observe that both SFT and SICL yield better performance and lower ECE with more training data, whereas ICL maintains similar performance. This indicates that supervised methods can better calibrate the model, and produce more calibrated predictions if provided with sufficient data. We also find that self-ensembling remains effective in improving model’s performance and mitigating the calibration error on average.
The findings hold across different model sizes. In order to assess the impact of model sizes, we further conduct experiments with Flan-T5xl. For full detailed experiment results, we refer the readers to Table 13 and 14 in the Appendix. We find that a larger model achieves a better calibration score with SFT and SICL than ZSL and ICL. We also witness similar behavior of ICL, surpassing SFT and SICL with or without Batch Calibration on seen data, which aligns with previous findings. The results on SST-5 and Hate Speech show that by applying the self-ensembling method, Flan-T5xl achieves better performance and lower calibration scores, indicating that the model becomes more ‘task-specialized’. It is also worth noticing that our method is able to improve the performance and decrease the calibration scores on some tasks (Hate Speech) where traditional calibration currently fails.
@@ 7 Conclusion
We have provided a comprehensive analysis of the intricate relationship between in-task performance and calibration across various learning methods in low-resource scenarios. Our findings illuminate the nuances of in-task performance and calibration across different task families, meanwhile addressing the inherent miscalibration over all learning methods. We have also investigated effective strategies to enhance both aspects simultaneously, offering a viable solution through self-ensembling: it results in more calibrated predictions with comparable or superior task performance. We hope that this study will contribute valuable insights into the dynamic landscape of LMs. These discoveries also offer practical guidance to practitioners, aiding them in choosing suitable learning paradigms and paving the way for the development of more reliable and high-performing LMs across diverse applications.
@@
Our experimental results are conducted with the Flan-T5 model family, which is an encoder-decoder architecture, where we have not investigated the behaviour of other popular choices of decoder-only models, such as Llama Touvron et al. (2023) in low-resource scenarios, with different learning methods and self-ensembling strategies. Secondly, limited by the training resources, our experiments only consider LMs within the 3B parameter budget. We will endeavor to scale our experiments to cover larger language models as part of future work. Moreover, there is a variety of additional, more sophisticated ensembling methods Mohammed and Kora (2023), where we have only studied the max, mean, and majority vote variants for self-ensembling. In future work, we aim to extend the analysis and the self-ensembling methods to more families of tasks, diverse types of models, and other ensembling techniques.
@@ B.2 Results of self-ensembling
Appendix C Supplementary Results
C.1 Comparison with Classical Calibration Methods
In comparison to self-ensembling in our work, both temperature scaling and Platt scaling Guo et al. (2017) require supervised labeled data as a validation set and parameter tuning (e.g., temperature) for different instruction templates, ICL settings, and models. In contrast, self-ensembling is tuning-free and employs strategies that are universally applicable across different learning methods. Furthermore, temperature scaling merely adjusts confidence levels while preserving the ranking among choices in classification tasks, making it ineffective in enhancing performance, as shown in Table 11. Platt scaling is mostly suitable for binary classification tasks, making it less applicable to multi-class classification tasks.
We argue that self-ensembling is orthogonal to other calibration methods, allowing other calibration methods to be applied on top of the ensembled results for further calibration with improved performance. It can help mitigate the problem of miscalibration, offering new perspectives on calibration in addition to previous methods and enhancing trustworthiness.
Through experiments, we also observe that temperature scaling is sensitive to the selection of supervised labelled data. If we choose a batch of label-imbalanced supervised labelled data, temperature scaling may fail in this case. On the other hand, self-ensembling method doesn’t have this problem due to its training-free nature, as shown in Table 12.
SST-5
ZSL ICL SFT SICL
Performance ECE Performance ECE Performance ECE Performance ECE
Original* 38.08 0.2305 37.88 0.1802 50.28 0.4307 44.27 0.3931
Temperature Scaling* 38.08 0.0626 37.88 0.0693 50.28 0.0804 44.27 0.0783
Original 42.00 0.1416 37.59 0.1833 46.27 0.403 47.12 0.3602
Self-Ensembling 46.79 0.0782 37.75 0.0955 48.35 0.3008 47.91 0.1200
Hate Speech
ZSL ICL SFT SICL
Performance ECE Performance ECE Performance ECE Performance ECE
Original* 47.08 0.2327 40.21 0.2729 57.41 0.3952 57.50 0.2526
Temperature Scaling* 47.08 0.0658 40.21 0.0684 57.41 0.0539 57.50 0.0312
Original 37.08 0.3175 40.09 0.2708 58.01 0.3541 59.48 0.1928
Self-Ensembling 36.95 0.2698 41.19 0.1596 57.19 0.2386 60.22 0.1151
Manifestos
ZSL ICL SFT SICL
Performance ECE Performance ECE Performance ECE Performance ECE
Original* 14.41 0.4362 19.33 0.4741 37.27 0.2591 35.87 0.2115
Temperature Scaling* 14.41 0.0781 19.33 0.0800 37.27 0.0630 35.87 0.0930
Original 14.50 0.4319 13.01 0.4760 35.76 0.2005 37.55 0.2138
Self-Ensembling 15.29 0.2620 13.69 0.1932 36.49 0.0662 37.61 0.1236
Table 11: Comparison with Temperature Scaling and Self-Ensembling. Self-ensembling helps to improve both performance and calibration where temperature scaling falls short in improving performance, making it a different perspective in calibration compared to temperature scaling. * indicates that the experiment results are only reported with 1 random seed.
ZSL ICL SFT SICL
Performance ECE Performance ECE Performance ECE Performance ECE
Original* 14.41 0.4362 19.33 0.4741 37.27 0.2591 35.87 0.2115
Temperature Scaling* 14.41 0.9333 19.33 0.9333 37.27 0.9333 35.87 0.9333
Original 14.50 0.4319 13.01 0.4760 35.76 0.2005 37.55 0.2138
Self-Ensembling 15.29 0.2620 13.69 0.1932 36.49 0.0662 37.61 0.1236
Table 12: Temperature scaling is sensitive to the selection of supervised labelled data. When choosing a batch of label-imbalanced supervised labelled data from Manifestos, temperature scaling may fail in this case. On the other hand, self-ensembling method doesn’t have this problem due to its training-free nature. * indicates that the experiment results are only reported with 1 random seed.
C.2 How about larger models?